OpenStudy (christos):

Factor 15x^2 - 37x + 20

hartnn (hartnn):

do you know how to generally factor a quadratic ?

OpenStudy (christos):

I know some methods but I guess I miss this one

hartnn (hartnn):

for \(ax^2+bx+c\), you need to find 2 numbers with sum = b and product = ac so, for your question, you need to find 2 numbers with sum =-37 and product = 20*15 = 300

OpenStudy (christos):

I see now :)

hartnn (hartnn):

so which 2 numbers u get ? then split -37x accordingly...

OpenStudy (christos):

I got -20 and -15 what now?

hartnn (hartnn):

good! split -37x as -20x-15x so, you have \(15x^2-15x-20x+20\) there's a common factor in 1st 2 terms and a common factor in last 2 terms, can you proceed ?

hartnn (hartnn):

oh wait, -20 and -15 are incorrect :P

OpenStudy (christos):

Yea :S

hartnn (hartnn):

try another pair ?

OpenStudy (christos):

-12 -25 ?

hartnn (hartnn):

bingo! split -37x as -25x-12x so, you have \(15x^2−25x−12x+20\) there's a common factor in 1st 2 terms and a common factor in last 2 terms, can you proceed ?

OpenStudy (christos):

x(x15-25)4(3x + 5)

hartnn (hartnn):

from first 2 terms factor out 5x

OpenStudy (christos):

5x(3x - 5) - 4(3x + 20)

hartnn (hartnn):

-12x+20 = -4 (3x+.....?)

OpenStudy (christos):

its 5 sorry

hartnn (hartnn):

+5 or -5 ?

OpenStudy (christos):

-5

hartnn (hartnn):

so, now factor 3x-5 from that

OpenStudy (christos):

What do you mean? I dont understand :SS

hartnn (hartnn):

5x (3x-5) -4(3x-5) factor out 3x-5 (3x-5)(5x-4) i meant this, and you're done

OpenStudy (anonymous):

Bonjour! Comment allez vous?

hartnn (hartnn):

i am fine, do u need any help ? with what ?

OpenStudy (anonymous):

You are very good at teaching people

hartnn (hartnn):

oh, thank you very much :) :D

OpenStudy (anonymous):

\[15x^2 - 37x + 20 \] \[(3 x-5) (5 x-4)\]